# Colourings of Cayley graphs of finite $3$-groups

### 20 Mar 2026 v1

@@ Abstract.
Colouring problems arising from group-based constructions provide a natural link between combinatorics and algebra, particularly in the study of Cayley graphs and Latin squares. We introduce colouring bijections of finite groups, a class of permutations encoding proper vertex colourings of associated Cayley-type graphs, extending classical notions such as complete and strong complete mappings. We prove that every finite $3$-group without a cyclic maximal subgroup admits a colouring bijection. Consequently, for such groups $G$, the graph $\mathscr{G}_{3}(G)$ admits a proper colouring with $|G|$ colours. These results show that the existence of colouring bijections is governed by structural properties of $3$-groups, revealing a new connection between group theory and combinatorial colouring problems.
@@
To study the equality case, we introduce the following notion. A bijection $\sigma\colon G\to G$ is called a colouring bijection if the maps
@@
are bijections of $G$. The existence of such a bijection implies that $\mathscr{G}_{3}(G)$ admits a proper colouring with $|G|$ colours (see Proposition 1.1). Thus, colouring bijections may be viewed as a natural higher-dimensional analogue of complete mappings in the context of graph colourings arising from Latin squares. A group admitting a colouring bijection will be called colourable.
In the abelian case, colouring bijections coincide with strong complete mappings, that is, bijections $\sigma\colon G\to G$ such that the maps $x\mapsto x\sigma(x)$ and $x\mapsto x^{-1}\sigma(x)$ are bijections. In general, there is no canonical transformation converting strong complete mappings into colouring bijections, since the expressions $\sigma(x)x$, $x^{-1}\sigma(x)$, and $x^{-1}\sigma(x)x$ are not related by uniform compositions with inversion, translations, or conjugation. The existence of strong complete mappings for abelian $3$-groups was determined by Evans [8], and further constructions were obtained in [4].
@@ 1. Preliminary results
([8]) Every noncyclic abelian $3$-group admits a strong complete mapping and hence is colourable.
@@
$\Delta^{(1)}_{\sigma}(ht)=(\psi(h)h)\,\varphi(t)t=(\psi(h)h\,\xi(t))\cdot t_{1},$
$\Delta^{(2)}_{\sigma}(ht)=(h^{-1}\psi(h))\,t^{-1}\varphi(t)=(h^{-1}\psi(h)\zeta(t))\cdot t_{2},$
$\Delta^{(3)}_{\sigma}(ht)=(h^{-1}\psi(h)h)\,t^{-1}\varphi(t)t=\psi(h)\omega(t)\cdot t_{3}.$
@@ 2. Colourability of groups of small order
We use coordinate notation for both groups. The permutations were obtained using the GAP system [13]. An exhaustive search was performed over permutations fixing the identity, and we verified whether the associated mapss $\Delta^{(1)}$, $\Delta^{(2)}$ and $\Delta^{(3)}$ were bijections.
@@ 3.1. Case $H=C_{3}\times C_{3}$
$\Delta^{(1)}_{\sigma}(ht)$ $=\sigma(ht)\cdot(ht)=\bigl(\psi_{t}(h)\,\varphi(t)\bigr)\cdot(ht)=\psi_{t}(h)\cdot\bigl(\varphi(t)\,h\,\varphi(t)^{-1}\bigr)\cdot\bigl(\varphi(t)\,t\bigr)$
$=\psi_{t}(h)\cdot\rho_{\varphi(t)}(h)\cdot\xi(t)\,t_{1}.$
$\Delta^{(2)}_{\sigma}(ht)$ $=(ht)^{-1}\sigma(ht)=t^{-1}h^{-1}\cdot\psi_{t}(h)\,\varphi(t)=t^{-1}\cdot\bigl(h^{-1}\psi_{t}(h)\bigr)\cdot\varphi(t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot(t^{-1}\varphi(t))=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\zeta(t)\,t_{2}.$
$\Delta^{(3)}_{\sigma}(ht)$ $=\Delta^{(2)}_{\sigma}(ht)\cdot(ht)=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot(t^{-1}\varphi(t))\cdot(ht)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\rho_{t^{-1}\varphi(t)}(h)\cdot(t^{-1}\varphi(t)t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\cdot\rho_{t^{-1}\varphi(t)}(h)\cdot\omega(t)t_{3}.$
@@
are invertible, and matrices of the form $C(\cdot)$ are invertible. Hence all $F_{i}^{(t)}$ are bijections $H\to H$.
@@
3.2. Case $H=C_{3}\times C_{9}$
We prove the existence of a lift of a colouring bijection of the $3$-group $G/H$ to a colouring bijection of $G$ in the case $H\cong C_{3}\times C_{9}$. The situation is asymmetric, since an element of order $9$ may or may not lie in the centre of $G$. This affects the form of the conjugations $\rho_{x}$ acting on $H$, and therefore the proof requires a separate treatment of these two cases.
@@ 3.2. Case $H=C_{3}\times C_{9}$
Let $G$ be a finite $3$-group and let $H\cong C_{3}\times C_{9}$ be an abelian normal subgroup such that the quotient group $G/H$ is colourable. Then $G$ is also colourable.
@@
Since $C_{3}\times C_{9}$ is colourable (see Theorem 1.3), by Lemma 1.6 it suffices to consider the case where $H$ is not contained in the centre of $G$.
@@
$\Delta^{(1)}_{\sigma}(ht)=\bigl(\alpha_{u(t)}(h)\,z^{u(t)j}h\bigr)(\varphi(t)t)=T_{u(t)}(h)(\varphi(t)t)=T_{u(t)}(h)\xi(t)\cdot t_{1}$
@@
$\Delta^{(2)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,(t^{-1}\varphi(t))=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,\zeta(t)\cdot t_{2},$
@@
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,(t^{-1}\varphi(t)t)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,\omega(t)\cdot t_{3}.$
@@
$\Delta^{(1)}_{\sigma}(ht)=A_{\lambda(t),\ell(t)}(h)\,\xi(t)\cdot t_{1}.$
@@
$=\rho_{t^{-1}}\!\bigl(B_{\lambda(t),\ell(t)}(h)\bigr)\,\zeta(t)\cdot t_{2}.$
@@
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(C_{\lambda(t),\ell(t)}(h)\bigr)\,\omega(t)\cdot t_{3}.$
@@ 4. Main theorem
    The colourability of the groups $L_{r}$ for $r\geq 4$ remains open. On the other hand, if $G$ is any nontrivial $3$-group, then $L_{r}\times G$ is colourable, since its centre is noncyclic and the product is therefore not isomorphic to any $L_{s}$.
@@
    By Theorem 4.3, every $3$-group without a cyclic subgroup of index $3$ is colourable. A natural next problem is the case of $2$-groups. It is plausible that the absence of cyclic subgroups of index $4$ is sufficient; the recent results of [2] on strong complete mappings support this expectation. Such a theorem would settle the colourability problem for finite nilpotent groups.
@@
  5. (5)
    For nonabelian groups, the three defining bijectivity conditions are genuinely independent: there exist bijections $\sigma$ for which two of the maps $\Delta^{(i)}_{\sigma}$ are bijective, while the third is not.
    In particular, for the group $L_{3}$ there exists a permutation $\sigma$ such that $\Delta^{(1)}_{\sigma}$ and $\Delta^{(2)}_{\sigma}$ are bijections, whereas $\Delta^{(3)}_{\sigma}$ is not. It is given by
    $\sigma=$ $(1\;ab\;a^{2}b^{2}\;a^{3}b\;a^{2}b\;a^{7}\;a^{7}b^{2}\;a\;a^{8}b^{2}\;a^{5}b\;a^{8}\;a^{6}\;a^{6}b\;a^{5}\;ab^{2}\;a^{2}\;a^{8}b\;a^{4}b^{2}\;a^{3})$
    $(b\;a^{4}b\;a^{3}b^{2}\;a^{5}b^{2}\;a^{6}b^{2}\;b^{2}\;a^{7}b),$
    where $a^{4}$ is a fixed point. This shows that, in the nonabelian setting, the three bijectivity conditions in the definition are not redundant.

### 24 Mar 2026 v2

@@ Abstract.
Colouring problems arising from group-based constructions provide a natural link between combinatorics and algebra, particularly in the study of Cayley graphs and Latin squares. We introduce the notion of colouring bijections of finite groups, a class of permutations encoding proper vertex colourings of associated Cayley-type graphs, extending classical concepts such as complete mappings and strong complete mappings.
We prove that every finite $3$-group without a cyclic maximal subgroup admits a colouring bijection. Consequently, for such a group $G$, the graph $\mathscr{G}_{3}(G)$ - a three-dimensional analogue of a Latin square - admits a proper colouring with $|G|$ colours.
These results show that the existence of colouring bijections is governed by structural properties of $3$-groups, revealing a new connection between group theory and combinatorial colouring problems.
@@
To investigate the equality case, we introduce the following notion. A bijection $\sigma\colon G\to G$ is called a colouring bijection if the maps
@@
are bijections of $G$. The existence of such a bijection implies that $\mathscr{G}_{3}(G)$ admits a proper colouring with $|G|$ colours (see Proposition 1.1). Thus, colouring bijections can be viewed as a natural higher-dimensional analogue of complete mappings in the context of graph colourings arising from Latin squares. A group admitting a colouring bijection will be called colourable.
In the abelian case, colouring bijections coincide with strong complete mappings, that is, bijections $\sigma\colon G\to G$ such that the maps $x\mapsto x\,\sigma(x)$ and $x\mapsto x^{-1}\,\sigma(x)$ are bijections. In general, however, there is a canonical transformation that converts colouring bijections into strong complete mappings. More precisely, as shown in Proposition 1.2, if $\sigma\colon G\to G$ is a colouring bijection, then $\tau=\sigma^{-1}$ is a strong complete mapping satisfying an additional conjugacy condition, namely that the map
$x\mapsto\tau(x)^{-1}\,x\,\tau(x)$
is also a bijection. This constitutes a significant restriction. Computational results for groups of small order indicate that colouring bijections are much rarer. For instance, in the modular group of order $16$, $M_{16}=\langle a,b\mid a^{8}=b^{2}=1,\;bab^{-1}=a^{5}\rangle,$ colouring bijections constitute only $\approx 2.17\%$ of all strong complete mappings.
The existence of strong complete mappings for abelian $3$-groups was determined by Evans [8], and further constructions were obtained in [4].
@@ 1. Preliminary results
We relate colouring bijections to strong complete mappings.
Let $\sigma\colon G\to G$ be a bijection such that the maps
$x\mapsto\sigma(x)\,x\qquad\text{and}\qquad x\mapsto x^{-1}\,\sigma(x)$
are bijections. Put $\theta=\sigma^{-1}$ and define $\theta^{c}\colon G\to G$ by
$\theta^{c}(x)=(\theta(x))^{-1}\,x\,\theta(x).$
Then
$\theta^{+}(x)=x\,\theta(x)=x\sigma^{-1}(x),\qquad\theta^{-}(x)=x^{-1}\,\theta(x)=x^{-1}\sigma^{-1}(x).$
Substituting $x=\sigma(y)$ we obtain
$\theta^{+}(\sigma(y))=\sigma(y)\,y=\Delta^{(1)}_{\sigma}(y),\qquad\theta^{-}(\sigma(y))=(\sigma(y))^{-1}y=\Delta^{(2)}_{\sigma}(y)^{-1}.$
Hence $\theta^{+}$ and $\theta^{-}$ are bijections if and only if $\Delta^{(1)}_{\sigma}$ and $\Delta^{(2)}_{\sigma}$ are bijections.
Substituting $x=\sigma(y)$ gives
$\Delta^{(3)}_{\sigma}(\sigma(y))=\sigma(y)^{-1}y\sigma(y)=\theta(y)^{-1}y\theta(y)=\theta^{c}(y).$
Thus $\Delta^{(3)}_{\sigma}$ is bijective if and only if $\theta^{c}$ is bijective.
Groups that admit strong complete mappings are referred to as strongly admissible in the literature. We summarize the preceding discussion as follows.
Proposition 1.2.
Let $G$ be a finite group and let $\sigma\colon G\to G$ be a bijection. Then
  1. (1)
    $\Delta^{(1)}_{\sigma}$ and $\Delta^{(2)}_{\sigma}$ are bijections if and only if $\sigma^{-1}$ is a strong complete mapping;
  2. (2)
    if $G$ is colourable, then $G$ is strongly admissible.
@@
([8]) Every noncyclic abelian $3$-group admits a strong complete mapping (consequently, is colourable).
@@
$\Delta^{(1)}_{\sigma}(ht)=(\psi(h)h)\,\varphi(t)t=(\psi(h)h\,\xi(t))\,t_{1},$
$\Delta^{(2)}_{\sigma}(ht)=(h^{-1}\psi(h))\,t^{-1}\varphi(t)=(h^{-1}\psi(h)\zeta(t))\,t_{2},$
$\Delta^{(3)}_{\sigma}(ht)=(h^{-1}\psi(h)h)\,t^{-1}\varphi(t)t=\psi(h)\omega(t)\,t_{3}.$
@@ 2. Colourability of groups of small order
We use coordinate notation for both groups. The permutations were obtained using the GAP system [13]. An exhaustive search was performed over permutations fixing the identity, and we verified whether the associated maps $\Delta^{(1)}$, $\Delta^{(2)}$ and $\Delta^{(3)}$ were bijections.
@@ 3.1. Case $H=C_{3}\times C_{3}$
$\Delta^{(1)}_{\sigma}(ht)$ $=\sigma(ht)\,(ht)=\bigl(\psi_{t}(h)\,\varphi(t)\bigr)\,(ht)=\psi_{t}(h)\,\bigl(\varphi(t)\,h\,\varphi(t)^{-1}\bigr)\,\bigl(\varphi(t)\,t\bigr)$
$=\psi_{t}(h)\,\rho_{\varphi(t)}(h)\,\xi(t)\,t_{1}.$
$\Delta^{(2)}_{\sigma}(ht)$ $=(ht)^{-1}\sigma(ht)=t^{-1}h^{-1}\,\psi_{t}(h)\,\varphi(t)=t^{-1}\,\bigl(h^{-1}\psi_{t}(h)\bigr)\,\varphi(t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\,(t^{-1}\varphi(t))=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\,\zeta(t)\,t_{2}.$
$\Delta^{(3)}_{\sigma}(ht)$ $=\Delta^{(2)}_{\sigma}(ht)\,(ht)=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\,(t^{-1}\varphi(t))\,(ht)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\,\rho_{t^{-1}\varphi(t)}(h)\,(t^{-1}\varphi(t)t)$
$=\rho_{t^{-1}}\Bigl(h^{-1}\psi_{t}(h)\Bigr)\,\rho_{t^{-1}\varphi(t)}(h)\,\omega(t)t_{3}.$
@@
are invertible, and matrices of the form C()C(\,) are invertible. Hence all $F_{i}^{(t)}$ are bijections $H\to H$.
@@
3.2. Case $H=C_{9}\times C_{3}$
We prove the existence of a lift of a colouring bijection of the $3$-group $G/H$ to a colouring bijection of $G$ in the case $H\cong C_{9}\times C_{3}$. The situation is asymmetric, since an element of order $9$ may or may not lie in the centre of $G$. This affects the form of the conjugations $\rho_{x}$ acting on $H$, and therefore the proof requires a separate treatment of these two cases.
@@ 3.2. Case $H=C_{9}\times C_{3}$
Let $G$ be a finite $3$-group and let $H\cong C_{9}\times C_{3}$ be an abelian normal subgroup such that the quotient group $G/H$ is colourable. Then $G$ is also colourable.
@@
Since $C_{9}\times C_{3}$ is colourable (see Theorem 1.4), by Lemma 1.7 it suffices to consider the case where $H$ is not contained in the centre of $G$.
@@
$\Delta^{(1)}_{\sigma}(ht)=\bigl(\alpha_{u(t)}(h)\,z^{u(t)j}h\bigr)(\varphi(t)t)=T_{u(t)}(h)(\varphi(t)t)=T_{u(t)}(h)\xi(t)\,t_{1}$
@@
$\Delta^{(2)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,(t^{-1}\varphi(t))=\rho_{t^{-1}}\!\bigl(\Delta^{(2)}_{\alpha_{u(t)}}(h)\bigr)\,\zeta(t)\,t_{2},$
@@
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,(t^{-1}\varphi(t)t)=\rho_{t^{-1}}\!\bigl(T_{u(t)}(h)\bigr)\,\omega(t)\,t_{3}.$
@@
$\Delta^{(1)}_{\sigma}(ht)=A_{\lambda(t),\ell(t)}(h)\,\xi(t)\,t_{1}.$
@@
$=\rho_{t^{-1}}\!\bigl(B_{\lambda(t),\ell(t)}(h)\bigr)\,\zeta(t)\,t_{2}.$
@@
$\Delta^{(3)}_{\sigma}(ht)=\rho_{t^{-1}}\!\bigl(C_{\lambda(t),\ell(t)}(h)\bigr)\,\omega(t)t_{3}.$
@@ 4. Main theorem
    The colourability of the groups $L_{r}$ for $r\geq 4$ remains open.
@@
    If $G$ and $H$ are nontrivial $3$-groups, then the direct product $G\times H$ is colourable. Indeed, its centre satisfies
    $Z(G\times H)=Z(G)\times Z(H),$
    which is noncyclic as a direct product of nontrivial abelian $3$-groups. Consequently, $G\times H$ is not isomorphic to any group $L_{s}$.
  4. (4)
    By Theorem 4.3, every $3$-group without a cyclic subgroup of index $3$ is colourable. A natural direction for further study is the case of $2$-groups. Recent results of [2] on strong complete mappings for $2$-groups suggest that an analogous phenomenon may occur in this setting. In particular, it is of interest to determine which strongly admissible $2$-groups are colourable.
    It is therefore natural to investigate colourability within the following families of $2$-groups: the dihedral groups $D_{2^{n}}$, the semidihedral groups $SD_{2^{n}}$, the generalized quaternion groups $Q_{2^{n}}$, and the modular groups $M_{2^{n}}$. These families comprise the principal classes of non-abelian $2$-groups of maximal class.